Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2000-211_2000-217

Mission Operations Days: 2000/211 to 2000/217
July 29, 2000 00:00:00 GMT to August 04, 2000 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - A Side
MODIS has an anomaly with the LWIR focal plane. See description below.
ComponentState(s)Comment
BlackbodyA On (315 K); B OffNominal
Calibration ElectronicsA On; B OffNominal
Control ProcessorA On; B OffNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA On; B OffNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 On; #3 & #4 OffNominal
Format ProcessorA On; B OffNominal
Power Supplies#1 On, #2 OffNominal
PV VIS FPAA On; B OffNominal
PV NIR FPAA On; B OffNominal
PV SMIR FPAA On; B OffNominal
PV LWIR FPAA On; B OffNominal
PC LWIR FPAA On; B OffNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (85 K)CFPA Anomaly Investigation
RC SMIR FPA HeaterOffNominal
Scan AssemblyA On; B OffNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA On; B OffNominal
Flight SoftwareRev BD + 3 sets of patches
  • FR Band 24 Gain Change (10 commands)
  • FR Band 25 Gain Change (10 commands)
  • FR Band 27 Gain Change (10 commands)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled66, 67Nominal

This Week’s Completed MODIS Activities:

Saturday, July 29, 2000

None

Sunday, July 30, 2000

None

Monday, July 31, 2000

2000/213 00:16:47 ATC – PV Ecal
2000/213 00:20:47 ATC – PC Ecal
2000/213 12:26:06 – 12:59:39 ATC - OA19 SRCA Full Radiometric
2000/213 12:00:00 ATC - Blackbody to 270K
2000/213 13:28:05 – 13:37:09 Data Loss

Tuesday, August 1, 2000

2000/214 06:00:00 ATC Blackbody to 280K
2000/214 10:00:00 ATC Blackbody to 285K
2000/214 12:00:00 ATC Blackbody to 290K
2000/214 14:00:00 ATC Blackbody to 295K
2000/214 14:00:30 ATC Set Blackbody duty cycle to 100%
2000/214 16:00:00 ATC Blackbody to 300K
2000/214 18:00:00 ATC Blackbody to 315K
2000/214 20:05:00 ATC Blackbody to 270K
2000/214 20:05:30 ATC Set Blackbody duty cycle to 33%

Wednesday, August 2, 2000

2000/215 12:13:43 – 13:11:14 ATC – OA23 SRCA Full Spatial
2000/215 19:00:00 ATC Blackbody to 290K
2000/215 19:30 Drag Make-up Maneuver

Thursday, August 3, 2000

2000/216 05:29:43 ATC OA15 - SD/SDSM Open
2000/216 07:08:36 ATC OA16 - SD/SDSM Screened
2000/216 07:57:25 – 08:49:04 ATC – OA22 SRCA Full Spectral, 30W part I
2000/216 09:52:13 – 10:26:58 ATC – OA22 SRCA Full Spectral, 30W part II
2000/216 11:14:41 – 12:05:57 ATC – OA22 SRCA Full Spectral, 10W part I
2000/216 13:10:00 –14:01:16 ATC – OA22 SRCA Full Spectral, 10W part II
2000/216 20:02:00 ATC Blackbody to 270K
2000/216 20:04:12 Real-time – Turn on LW FPA Heater
2000/216 20:04:59 Real-time – Set focal plane temp to 85
2000/216 20:48 - 21:01 Real-time – Set new gains for bands 24, 25, 27
2000/216 21:35 Real-time – Dump Gain table failed due to timing issues
2000/216 22:24:55 Real-time – Dump Gain Table successful

Friday, August 4, 2000

2000/217 17:00:00 ATC Blackbody to 315K
2000/217 17:00:30 ATC Blackbody duty cycle to 100%

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Saturday, August 5, 2000

2000/218 00:00:00 ATC Blackbody to 270K
2000/218 00:00:30 ATC Blackbody duty cycle to 33%
2000/218 22:00:00 ATC Blackbody to 315K
2000/218 22:00:30 ATC Blackbody duty cycle to 100%

Sunday, August 6, 2000

2000/219 05:00:00 ATC Blackbody to 270K
2000/219 05:00:30 ATC Blackbody duty cycle to 33%

Monday, August 7, 2000

2000/220 03:00:00 ATC Blackbody to 290K
2000/220 14:00:?? Real-time – Set Itwk/Vdet to 79/190
2000/220 14:00:?? Real-time – Transition to Standby
2000/220 14:00:?? Real-time – Transition to Low Power
2000/220 14:00:?? Real-time – Turn on Cold Stage
2000/220 14:30:?? Real-time – Turn on Intermediate Stage
2000/220 16:30:00:?? Real-time – Turn on Outer Stage
2000/220 19:00 Attain 318K

Tuesday, August 8, 2000

2000/221 19:00:?? Real-time – Transition to Standby (timing depends on heater temperature warm-up rate)
2000/221 23:00:?? Real-time – Transition to Science
2000/221 23:00:?? Real-time – Turn on Blackbody
2000/221 23:00:?? Real-time – Set Blackbody to 285K

Wednesday, August 9, 2000

None

Thursday, August 10, 2000

None
2000/223 23:00 Should be controlling at 83K today

Friday, August 11, 2000

None

Maneuvers:

Wednesday, August 2, 2000 - Drag makeup maneuver (215/19:30:00)

MODIS Anomalies:

The Radiative Cooler has negative margin, therefore the cold focal planes are no longer under thermal control. This affects bands 5, 6, 7, and 20-36. Analysis is underway.

It appears that the gains for band 14H were reset to the original launch values some time between 2000/173/23:45 – 2000/174/03:55. Investigation is underway to determine the reason.

General Instrument Comments:

MODIS is in Science Mode on the A-side with the SVD and NAD open.

Gain Changes

BandDetectorCurrent GainNew Gain
241108102
24210094
2439792
2449893
24510296
24610195
2479489
2489388
2499994
24109489
251148136
252127127
253120114
254115109
255119113
256115109
257119113
258109103
259109103
2510115109
271238228
272219209
273231221
274238238
275238238
276238238
277238238
278238238
279238238
2710238238

MODIS Telemetry Trends:

See MODIS Anomalies section.

Non-MODIS Significant Events:

Wednesday, July 19 (Day 201) the EOS Operations Center switched to the upgraded on-line operations software build, Rev C. They will also upgraded to Project Database (PDB) 28. The switchover went well and the MODIS IOT is comfortable with the upgrades.

Data processing by EDOS is currently running at a "degraded level". For a detailed summary of the events that have affected their products, view the EOS System Status web page (http://jupiter02.gsfc.nasa.gov:591/sysstat/) and select the EDOS summary link at the top of the page. A key item to note is from day 6/27/2000 at 12:11:19 PM: " The second major issue is the amount of data that EDOS is receiving since ASTER has been fully turned on." With the new, higher data output of ASTER, even less "spare" time will be available to EDOS for the processing of backlogged or reprocessing of data.

Limited Life Item Status:

SRCA 10W Lamp #1: 173.3 of 500 hours
SRCA 10W Lamp #2: 135.9 of 500 hours
SRCA 10W Lamp #3: 145.1 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 555.8 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1393 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements